# Ravensburg

### v3

@@ 3.1 Technical details
GenL has been developed in the high-level programming language MATLAB (The MathWorks Inc., Natick, MA, USA). MATLAB is available for numerous operating systems, namely Windows, macOS, and Linux/Unix. The GenL program can either run from the command line after adjusting the input file in a text editor or the GenL GUI can be used. Both methods are equivalent. Since the code is rather compact, an adaptation to other programming languages such as Python, C, C++, or Java is straightforward. However, any alteration of the GUI requires a deeper understanding of the MATLAB programming language. The program was tested on the following MATLAB versions: R2019b, R2020b, R2022b, R2023a, and R2023b. No additional toolbox installations are required. The GenL GUI runs correctly in MATLAB versions published in 2022 or later. In older versions, the graphical display in the app might be distorted. Possible distortions of the graphical display in all versions can be avoided if the program is running on one screen only.
@@ 3.2 Using GenL
For running GenL from the command line, the user is required to edit the input file ”fit_laue_oscillations.m”. This input file is displayed as a green trapeze shaped box in the flow chart in Fig. 2. First, file and path of the measurement as well as important measurement parameters need to be specified. Second, material specific parameters (atomic number , density, chemical composition, the Debye-Waller pre-factor), sample parameters (layer thickness, expected Bragg angle/interplanar spacing), and instrument parameters (wavelength, setup, measurement time, amplitude, resolution, and background level) need to be added. For a limited number of elements, the element-specific Debye-Waller pre-factor can be extracted from a database [Sears1991] as well or be put in manually. Furthermore, three features can be included in the fitting process of the program: strain, roughness, and a potentially present peak originating from a single-crystalline substrate can be accounted for by adding a Lorentzian intensity contribution at the respective Bragg angle. If any of these features are included, their respective parameters need to be specified in the code as well. Finally, for all fitting parameters upper and lower bounds are constructed and can be adjusted as necessary. Running the input file will start the fitting process, which is further described in the next section.
@@ 4 Application examples
In this section we showcase some examples of employing GenL to fit diffraction patterns measured on W and Fe thin films. Details on the thin film growth can be found in the respective publications for the W [Ravensburg2023] and the Fe thin films [Ravensburg2022, Ravensburg2024Fe]. All diffractograms presented in this section were measured in a Bede D1 diffractometer equipped with a Cu x-ray source operated at 35 mA and 50 kV. A circular mask (diameter: 0.005 m) and an incidence and a detector slit (both 0.0005 m) were used. The beam was monochromatized by reducing the CuK and CuK radiation using a Göbel mirror and a 2-bounce-crystal on the incidence side. The x-rays were detected with a Bede EDRc x-ray detector. Details on growth and diffraction study of an Fe/V superlattice with a ratio of 4 monolayers of Fe to 28 monolayers of V in 11 bilayer repetitions, which is presented in comparison to a calculation with GenL, can be found elsewhere [Droulias2017].
@@ 4.1 The influence of roughness
Thin W () layers can be epitaxially grown on AlO () substrates with high crystal quality despite a large lattice mismatch of 7.2 % and 19.4 % along W [] and W [] directions [Ravensburg2023]. The diffraction pattern of such a thin film, as shown in Fig. 4, exhibits Laue oscillations around the W () Bragg peak. The Laue oscillations have a high degree of symmetry indicating a small degree of strain in the layer. Fitting the oscillations in GenL yields the fit displayed in blue in Fig. 4. The fit captures the main Bragg peak position and the spacing of the oscillations well. However, the ratio of the height of the oscillations in relation to the main Bragg peak height is overestimated. Taking interface roughness into account in GenL, yields a better match of the relative intensity ratio. The corresponding fit estimating the roughness to be 4 Å is displayed in red. This interface roughness obtained from the GenL fit is in agreement with roughness values at the W/capping interface obtained from Kiessig fringe fitting using GenX [Bjorck2007, Glavic_Bjorck_2022] of the same sample [Ravensburg2023]. The analysis of the GenL fitting employed in [Ravensburg2023] contributed to the understanding that W exhibits a semicoherent interface on AlO () with an immediate strain relaxation taking place directly at the substrate/W interface and serves as an example of the successful application of GenL in studies on epitaxial growth.
@@ 4.2 Accounting for overlapping substrate peak intensity
A similar analysis can also be done with the scattering pattern around the W () Bragg peak instead, which is displayed in Fig. 5. However, the sharp peak at around 38 degrees originating from the single crystalline AlO () substrate overlaps in intensity with the Laue oscillations, making an analysis more challenging. This angular overlap in the scattering patterns originating from film and substrate is common for epitaxial growth as it is attributed to the similar out-of-plane lattice parameters in the substrate and the film. In GenL, the substrate peak intensity can be modeled with a Lorentzian profile, which is added to the scattering intensity attributed to the film. An example of a fit including roughness together with this modeled substrate peak intensity is displayed in red. Both fits from Fig. 4 and Fig. 5 for the same sample, yield similar fitted parameters. The deviation of the obtained from both fits is 0.05 %.
@@ 4.3 The influence of strain
Besides roughness and substrate peak intensity, GenL is also capable of taking a strain profile, i.e., a change in out-of-plane interplanar spacing over layer thickness into account, similar to [Lichtensteiger2018]. In Fig. 6 the diffraction pattern of a 100 Å thick Fe layer grown on an MgAlO () substrate is displayed. The GenL fit, employing an exponentially decaying strain profile, is shown in blue. Based on the fit, the out-of-plane lattice spacing decreases over Fe layer thickness, in line with a tensile out-of-plane strain induced at the substrate/Fe interface by a lattice mismatch [Ravensburg2022]. Moreover, the shape of this strain profile is in agreement with the observation that thinner Fe films, which are epitaxially grown on MgAlO (), exhibit on average larger out-of-plane lattice spacings in the order of the predicted values in the profile. However, the fit does not fully capture the asymmetric decay of the oscillations on both sides of the Bragg peak. In an experimental study on thin Fe layers (6 to 100 Å) grown on MgAlO () substrates [Ravensburg2024Fe], Fe grows in a tetragonally distorted, body centered tetragonal (bct) crystal structure for Fe layer thicknesses below 3 to 8 ML before the equilibrium body centered cubic (bcc) crystal structure stabilizes on this substrate. GenL can be used to investigate this growth even at the thin film limit. Hence, the diffraction pattern was fitted employing the out-of-plane interplanar distance profile of a bct Fe/bcc Fe bilayer with both bilayer thicknesses and interplanar spacings as fitting parameters. The results are displayed in red in Fig. 6. It is evident that the fit based on the bilayer profile captures the detected scattering intensity better and reproduces the asymmetry present in the intensity of the oscillations around the Fe () Bragg peak. This example demonstrates the full potential of the application of GenL in research work.
@@ 4.4 Limitations and possible extensions of the program
Besides the factors already included in GenL, there are others that may also influence the measured diffracted intensity from an epitaxial thin film, e.g., sample alignment, the x-ray beam footprint, crystallographic defects, or mosaicity, all of which are not yet included in the program. Some of these factors will be added in future versions of GenL. Currently, epitaxial thin films composed of more than one element (alloys) can only be modeled via a composition and an average density over the layer, not fully accounting for the correct atom stacking in the unit cell. Including correct atom stacking in alloys requires the computation of a structure factor for the specific crystallographic positions of the respective atoms in the unit cell. Details on such calculations can be found elsewhere [Lichtensteiger2018]. However, the command line version of GenL can be extended to correctly calculate diffraction patterns for these alloys due to the fact that GenL is modeling an effective one-dimensional stack of atoms whose order and positions can be adjusted. Even different/additional layers with multiple repetitions can be included. Hence, GenL can be easily adapted for more complex multilayer stacks composed of different materials and/or different crystal structures, such as superlattice structures. An example of a diffraction pattern from an Fe/V superlattice simulated in GenL, is displayed in Fig. 7 and compared against experimentally measured diffraction data [Droulias2017]. Fitting of superlattice diffraction data is not yet implemented in the current version of GenL, but will be included in future versions. The simulation of superlattice diffraction patterns is, however, straight forward. A version of the GUI, which can simulate superlattice diffraction patterns, is available in the program repository. Moreover, the code versatility also enables the user to conduct a more in-depth analysis by simulating the presence of atomic step terraces. In particular for epitaxial systems where out-of-plane atomic steps in the substrate are incommensurate with the out-of-plane atomic distance in the epitaxially growing film, interference of the scattering intensity from different stacks with vertical mismatch can contribute to an observed asymmetry in Laue oscillations [Fullerton1992, Miller2022], which may be relevant for fitting certain scattering patterns.
@@ 6 Program distribution
GenL is freely available from the authors under the GNU General Public Licence (GPL). A web page (URL: https://github.com/scatterer/GenL) exists for the distribution of the program. The package contains the Matlab code, an executable of the GUI, example files for fitting and examples for sample simulations if the GUI is used. Comments, feedback, and support can be communicated using the following email addresses: gunnar.palsson@physics.uu.se and vassilios.kapaklis@physics.uu.se.

### v4

@@ 4.1 Technical details
GenL has been developed in the high-level programming language MATLAB (The MathWorks Inc., Natick, MA, USA). MATLAB is available for numerous operating systems, namely Windows, macOS, and Linux/Unix. The GenL program can either run from the command line after adjusting the input file in a text editor, or the GenL GUI can be used. The GUI currently supports fitting a single layer using the kinematic approximation whereas the command line version supports any structure and either the kinematic or the dynamic approach. Since the code is rather compact, an adaptation to other programming languages such as Python, C, C++, or Java is straightforward. However, any alteration of the GUI requires a deeper understanding of the MATLAB programming language. The command line version of the program was tested on the following MATLAB versions: R2019b, R2020b, R2022b, R2023a, and R2023b. No additional toolbox installations are required. The GenL GUI runs correctly in MATLAB versions published in 2022 or later. In older versions, the graphical display in the app might be distorted. Possible distortions of the graphical display in all versions can be avoided if the program is running on one screen only.
@@ 4.2 Using GenL
For running GenL from the command line, the user is required to edit the input file ”fit_laue_oscillations.m”. This input file is displayed as a green trapeze shaped box in the flow chart in Fig. 4. First, file and path of the measurement as well as important measurement parameters need to be specified. Second, material specific parameters (atomic number , density, chemical composition, the Debye-Waller pre-factor), sample parameters (layer thickness, expected Bragg angle/interplanar spacing), and instrument parameters (wavelength, setup, measurement time, amplitude, resolution, and background level) need to be added. For a limited number of elements, the element-specific Debye-Waller pre-factor can be extracted from a database [Sears1991] as well, or be put in manually. Furthermore, three features can be included in the fitting process of the program: strain, roughness, and a potentially present peak originating from a single-crystalline substrate can be accounted for by adding a Lorentzian intensity contribution at the respective Bragg angle. If any of these features are included, their respective parameters need to be specified in the code as well. Finally, for all upper and lower bounds for the fitting parameters are constructed and can be adjusted as necessary. Running the input file will start the fitting process, which is further described in the next section.
@@ 5 Application examples
In this section we showcase some examples of employing GenL to fit diffraction patterns measured on W and Fe thin films. Details on the thin film growth can be found in the respective publications for the W [Ravensburg2023] and the Fe thin films [Ravensburg2022, bct_Fe_Anna]. All diffractograms presented in this section were measured in a Bede D1 diffractometer equipped with a Cu X-ray source operated at 35 mA and 50 kV. A circular mask (diameter: 0.005 m) and an incidence and a detector slit (both 0.0005 m) were used. The beam was monochromatized by reducing the CuK and CuK radiation using a Göbel mirror and a 2-bounce-crystal on the incidence side. The X-rays were detected with a Bede EDRc X-ray detector. Details on growth and diffraction study of a Fe/V superlattice with a ratio of 4 monolayers of Fe to 28 monolayers of V in 11 bilayer repetitions, which is presented in comparison to a calculation with GenL, can be found elsewhere [Droulias2017].
@@ 5.1 The influence of roughness
Thin W () layers can be epitaxially grown on Al2O3 () substrates with high crystal quality despite a large lattice mismatch of 7.2 % and 19.4 % along W [] and W [] directions [Ravensburg2023]. The diffraction pattern of such a thin film, as shown in Fig. 6, exhibits Laue oscillations around the W () Bragg peak. The Laue oscillations have a high degree of symmetry indicating a small degree of strain in the layer. Fitting the oscillations including roughness in GenL yields the fit displayed in red in Fig. 6. Based on the fit, the roughness is estimated to be 4 Å. This interface roughness is in agreement with roughness values at the W/capping interface obtained from Kiessig fringe fitting using GenX [Bjorck2007, Glavic_Bjorck_2022] of the same sample [Ravensburg2023]. The analysis of the GenL fitting employed in [Ravensburg2023] contributed to the understanding that W exhibits a semicoherent interface on Al2O3 () with an immediate strain relaxation taking place directly at the substrate/W interface and serves as an example of the successful application of GenL in studies on epitaxial growth.
@@ 5.2 Accounting for overlapping substrate peak intensity
A similar analysis can also be done with the scattering pattern around the W () Bragg peak instead, which is displayed in Fig. 7. However, the sharp peak at around 38 degrees originating from the single crystalline Al2O3 () substrate overlaps in intensity with the Laue oscillations, making an analysis more challenging. This angular overlap in the scattering patterns originating from film and substrate is common for epitaxial growth as it is attributed to the similar out-of-plane lattice parameters in the substrate and the film. An example of a fit including roughness together with this modeled substrate peak intensity is displayed in red. Both fits from Fig. 6 and Fig. 7 for the same sample, yield similar fitted parameters. The deviation of the obtained from both fits is 0.05 %.
@@ 5.3 The influence of strain
Besides roughness and substrate peak intensity, GenL is also capable of taking a strain profile, i.e., a change in out-of-plane interplanar spacing over layer thickness into account, similar to [Lichtensteiger2018]. In Fig. 8 the diffraction pattern of a 100 Å thick Fe layer grown on a MgAl2O4 () substrate is displayed. The GenL fit, employing an exponentially decaying strain profile, is shown in blue. In the employed strain profile, the out-of-plane lattice spacing decreases over Fe layer thickness, in line with a tensile out-of-plane strain induced at the substrate/Fe interface by a lattice mismatch [Ravensburg2022]. The shape of this strain profile is in agreement with the observation that thinner Fe films, which are epitaxially grown on MgAl2O4 (), exhibit on average larger out-of-plane lattice spacings in the order of the predicted values in the profile. However, the fit does not fully capture the asymmetric decay of the oscillations on both sides of the Bragg peak. In an experimental study on thin Fe layers (6 to 100 Å) grown on MgAl2O4 () substrates [bct_Fe_Anna], Fe grows in a tetragonally distorted, body centered tetragonal (bct) crystal structure for Fe layer thicknesses below 3 to 8 ML before the equilibrium body centered cubic (bcc) crystal structure stabilizes on this substrate. GenL can be used to investigate this growth even at the thin film limit. Hence, the diffraction pattern was fitted employing the out-of-plane interplanar distance profile of a bct Fe/bcc Fe bilayer with both bilayer thicknesses and interplanar spacings as fitting parameters [bct_Fe_Anna]. The results are displayed in red in Fig. 8. It is evident that the fit based on the bilayer profile captures the detected scattering intensity better and reproduces the asymmetry present in the intensity of the oscillations around the Fe () Bragg peak. This example demonstrates the full potential of the application of GenL in research work.
@@ 5.4 Limitations and possible extensions of the program
Besides the factors already included in GenL, there are others that may also influence the measured diffracted intensity from an epitaxial thin film, e.g., sample alignment, the X-ray beam footprint, crystallographic defects, or mosaicity, all of which are not yet included in the program. Some of these factors will be added in future versions of GenL.
The command line version of GenL can be extended to correctly calculate diffraction patterns for alloys due to the fact that GenL is modeling an effective one-dimensional stack of atoms whose order and positions can be adjusted. Even different/additional layers with multiple repetitions can be included. Hence, GenL can be easily adapted for more complex multilayer stacks composed of different materials and/or different crystal structures, such as superlattice structures. An example of a diffraction pattern from a Fe/V superlattice simulated in GenL, is displayed in Fig. 9 and compared against experimentally measured diffraction data [Droulias2017].
Fitting of superlattice diffraction data is not yet implemented in the current version of GenL, but will be included in future versions. The simulation of superlattice diffraction patterns is, however, straight forward in the command line version. Moreover, the code versatility also enables the user to conduct a more in-depth analysis by simulating the presence of atomic step terraces. In particular for epitaxial systems where out-of-plane atomic steps in the substrate are incommensurate with the out-of-plane atomic distance in the epitaxially grown film, interference of the scattering intensity from different stacks with vertical mismatch can contribute to an observed asymmetry in Laue oscillations [Fullerton1992, Miller2022], which may be relevant for fitting certain scattering patterns.
@@ 7 Program distribution
GenL is freely available from the authors under the GNU General Public Licence (GPL). A web page (URL: https://github.com/scatterer/GenL) exists for the distribution of the program. The package contains the Matlab code, an executable of the GUI, example files for fitting and examples for sample simulations if the GUI is used. Comments, feedback, and support can be communicated using the following email addresses: vassilios.kapaklis@physics.uu.se and gunnar.palsson@physics.uu.se.